# Efficient Inference for Large Language Model-based Generative Recommendation

### v2

@@ A.2 Proof and detailed derivatives
$=-\sum_{\bm{y}}q(\bm{y})\log\frac{1}{p(\bm{y})}-\sum_{\bm{y}}q(\bm{y})\log p(\bm{y}_{K})$
@@
According to the derivatives in Eq.(14), we obtain the alignment objective as shown in Eq.(2). Since we have $q(\bm{y})=\prod_{t}q(y_{t}|\bm{x},\bm{y}_{<t})=\prod_{t}q(y_{t}|c_{<t})$, we can further expand with some rearrangement into:
$-\sum_{\bm{y}}\prod_{t}q(y_{t}|c_{<t})\log\frac{\prod_{t}q(y_{t}|c_{<t})}{\prod_{i}p(y_{t}|c_{<t})}+\sum_{\bm{y}}\prod_{i}q(y_{t}|c_{<t})\log\frac{\prod_{i}q(y_{t}|c_{<t})}{\prod_{i}p(\bm{y}_{K,t}|\bm{y}_{K})}$ (15)
$=$ $-\sum_{\bm{y}\in\mathcal{Y}}\big{[}\frac{1}{|\bm{y}|}\sum_{t=1}^{|\bm{y}|}\big{[}{\sum_{y_{t}\in\mathcal{V}}q_{\theta}({y_{t}}|c_{<t})\log\frac{q_{\theta}(y_{t}|c_{<t})}{p(y_{t}|c_{<t})}}+{\sum_{y_{t}\in\mathcal{V}}q_{\theta}(y_{t}|c_{<t})\log\frac{q_{\theta}(y_{t}|c_{<t})}{p(y_{K})}}\big{]}\big{]},$
where $c_{t}=(\bm{x},\bm{y}_{<t})$ is the context, $p(y_{K})=p(\bm{y}_{K,t}|\bm{y}_{K,<t})$. Based on Eq.(15), we can easily obtain the alignment objective as in Eq.(3).
@@
Derivatives of Acceptance Rate for Relaxed Sampling Verification. Based on the with-replacement sampling approximation, for every independent sampling, we have acceptance rate of a sequence
@@
Based on the with-replacement sampling approximation, we can obtain the acceptance rate of a step with $K$ accepted sequences as $\beta=\prod_{K}b$. Similarly, we could also extend the distribution equivalence for sampling a sequence between the draft model and the target LLM from Appendix A.1 in  (Leviathan et al., 2023). As such, when the sampling-based beam search is performed with a beam size $K$, the joint distribution of the output $K$ sequences from the SD under the relaxed sampling verification is approximately equivalent to the counterparts from the target LLM, according to the approximation of with-replacement sampling, i.e., independent sampling.
@@ A.3 Experimental Details and Analysis
Table 3: Statistics of the two datasets.
@@
Utilization of Beam Sequences from Diverse Steps. Table 4 shows inference and training efficiency comparison between AtSpeed training with and without additional data generated from beam search results on diverse steps. From the table, it is observed that incorporating data from diverse generating steps usually enhances the speedup and accept steps, which is attributed to the additional alignment of top-$K$ sequences across different steps. Nevertheless, using additional data could incur increased training time and storage cost. Specifically, utilizing a larger dataset necessitates greater storage capacity and extends the training duration. Despite these additional costs, the marginal improvement in performance is not substantial, leading to a low cost-effectiveness ratio. Therefore, we should carefully consider the balance between the benefits of enhanced alignment against the increased training resource burden.
Hyper-parameter Analysis. 1) Effect of alignment strength $\alpha$. Figure 5 (b) and (c) presents the results indicating the impact of alignment strength on AtSpeed-R. As the value of $\alpha$ increases, the overall trend of accept steps exhibits an upward trajectory, which validates the necessity of alignment to improve the performance. 2) Effect of draft length $\gamma$. Figure 6 presents the impact of varying $\gamma$ on the performance of AtSpeed. Despite acceptance generally increases with $\gamma$, speedup does not consistently improve with $\gamma$. The trend of speedup in relation to $\gamma$ is influenced by draft beam size parameter $N$. Specifically, when $N$ is small, speedup tends to increase with $\gamma$; conversely, when $N$ is large, speedup tends to decrease with $\gamma$. This variation is primarily attributed to the computational overhead during the draft phase, which escalates with larger beam size, thereby diminishing the acceleration benefits. The key determinant of the draft phase’s computational cost is the constraint search process.

### v3

@@ A.2.2 Derivation of Alignment objectives of AtSpeed-S
$=-\sum_{\bm{y}}q(\bm{y})\log\frac{1}{p(\bm{y})}-\sum_{\bm{y}}q(\bm{y})\log p(\bm{y}_{K}).$
We can add $(\sum_{\bm{y}}q(\bm{y})\log q(\bm{y})-\sum_{\bm{y}}q(\bm{y})\log q(\bm{y}))$ to the expression since $\sum_{\bm{y}}q(\bm{y})\log q(\bm{y})-\sum_{\bm{y}}q(\bm{y})\log q(\bm{y})=0$. After adding the term with some rearrangement, we obtain
$-\sum_{\bm{y}}q(\bm{y})\log\frac{1}{p(\bm{y})}-\sum_{\bm{y}}q(\bm{y})\log p(\bm{y}_{K})+(\sum_{\bm{y}}q(\bm{y})\log q(\bm{y})-\sum_{\bm{y}}q(\bm{y})\log q(\bm{y}))$ (28)
$=-\sum_{\bm{y}}q(\bm{y})\log\frac{1}{p(\bm{y})}-\sum_{\bm{y}}q(\bm{y})\log q(\bm{y})+\sum_{\bm{y}}q(\bm{y})\log q(\bm{y})-\sum_{\bm{y}}q(\bm{y})\log p(\bm{y}_{K})$
@@
where $p(y_{K})=p(\bm{y}_{K,t}|\bm{y}_{K,<t})$. According to the derivation, we obtain the alignment objective as in Eq.(2). We now consider context in LLM-based recommendation. We omit input $\bm{x}$ and denote $c_{<t}=\bm{y}_{<t}$, referred to as context, then we have $q(\bm{y})=\prod_{t}q(y_{t}|c_{<t})$ and $p(\bm{y})=\prod_{t}p(y_{t}|c_{<t})$. We then can substitute $q(\bm{y})$ with $\prod_{t}q(y_{t}|c_{<t})$ and $p(\bm{y})$ with $\prod_{t}p(y_{t}|c_{<t})$ and rewrite the objective in Eq.(28) as:
$-\sum_{\bm{y}}q(\bm{y})\log\frac{q(\bm{y})}{p(\bm{y})}+\sum_{\bm{y}}q(\bm{y})\log\frac{q(\bm{y})}{p(\bm{y}_{K})}$ (29)
$=$ $-\sum_{\bm{y}}q(\bm{y})\log\frac{\prod_{t}q(y_{t}|c_{<t})}{\prod_{t}p(y_{t}|c_{<t})}+\sum_{\bm{y}}q(\bm{y})\log\frac{\prod_{t}q(y_{t}|c_{<t})}{\prod_{t}p(\bm{y}_{K,t}|\bm{y}_{K,<t})}$
$=$ $-\sum_{\bm{y}}q(\bm{y})[\log\prod_{t}q(y_{t}|c_{<t})-\log\prod_{t}p(y_{t}|c_{<t})]+\sum_{\bm{y}}q(\bm{y})[\log\prod_{t}q(y_{t}|c_{<t})-\log\prod_{t}p(\bm{y}_{K,t}|\bm{y}_{K,<t})]$
$=$ $-\sum_{\bm{y}}q(\bm{y})[\sum_{t}\log q(y_{t}|c_{<t})-\sum_{t}\log p(y_{t}|c_{<t})]+\sum_{\bm{y}}q(\bm{y})[\sum_{t}\log q(y_{t}|c_{<t})-\sum_{t}\log p(\bm{y}_{K,t}|\bm{y}_{K,<t})]$
$=$ $-\sum_{\bm{y}}q(\bm{y})\sum_{t}\log\frac{q(y_{t}|c_{<t})}{p(y_{t}|c_{<t})}+\sum_{\bm{y}}q(\bm{y})\sum_{t}\log\frac{q(y_{t}|c_{<t})}{p(\bm{y}_{K,t}|\bm{y}_{K,<t})}$
$=$ $-\sum_{t}\sum_{\bm{y}}q(\bm{y})\log\frac{q(y_{t}|c_{<t})}{p(y_{t}|c_{<t})}+\sum_{t}\sum_{\bm{y}}q(\bm{y})\log\frac{q(y_{t}|c_{<t})}{p(\bm{y}_{K,t}|\bm{y}_{K,<t})}.$
Since $\bm{y}=(y_{1},y_{2},\dots,y_{n})=(\bm{y}_{<t},y_{t},\bm{y}_{>t})$ is a sequence, we can decompose $\sum_{\bm{y}}q(\bm{y})$ into nested sum $\sum_{\bm{y}_{<t}}\sum_{y_{t}}\sum_{\bm{y}_{>t}}q(\bm{y})=\sum_{\bm{y}_{<t}}\sum_{y_{t}}\sum_{\bm{y}_{>t}}q(\bm{y}_{<t},y_{t},\bm{y}_{>t})$, where the nested sum over three parts of $\bm{y}$, i.e., $\bm{y}_{<t},\bm{y}_{t}$, and $\bm{y}_{>t}$ can cover all possible sequence $\bm{y}$. Since $q(\bm{y}_{<t},y_{t},\bm{y}_{>t})=q(\bm{y}_{<t})q(y_{t}|c_{<t})q(\bm{y}_{>t}|c_{\leq t})$, we can rewrite the nested sum of $q(\bm{y}_{<t},y_{t},\bm{y}_{>t})$ over $\bm{y}$:
$\sum_{\bm{y}}q(\bm{y})$ (30)
$=$ $\sum_{\bm{y}_{<t}}\sum_{y_{t}}\sum_{\bm{y}_{>t}}q(\bm{y}_{<t})q(y_{t}|c_{<t})q(\bm{y}_{>t}|c_{\leq t})$
$=$ $\sum_{\bm{y}_{<t}}\sum_{y_{t}}q(\bm{y}_{<t})q(y_{t}|c_{<t})\sum_{\bm{y}_{>t}}q(\bm{y}_{>t}|c_{\leq t})$
$=$ $\sum_{\bm{y}_{<t}}q(\bm{y}_{<t})\sum_{y_{t}}q(y_{t}|c_{<t})\sum_{\bm{y}_{>t}}q(\bm{y}_{>t}|c_{\leq t}).$
We then can substitute $\sum_{\bm{y}}q(\bm{y})$ with $\sum_{\bm{y}_{<t}}q(\bm{y}_{<t})\sum_{y_{t}}q(y_{t}|\bm{y}_{<t})\sum_{\bm{y}_{>t}}q(\bm{y}_{>t}|c_{\leq t})$ into Eq.(29) and obtain
$\quad-\sum_{t}\sum_{\bm{y}}q(\bm{y})\log\frac{q(y_{t}|c_{<t})}{p(y_{t}|c_{<t})}+\sum_{t}\sum_{\bm{y}}q(\bm{y})\log\frac{q(y_{t}|c_{<t})}{p(\bm{y}_{K,t}|\bm{y}_{K,<t})}$ (31)
$=-\sum_{t}[\sum_{\bm{y}_{<t}}q(\bm{y}_{<t})\sum_{y_{t}}q(y_{t}|c_{<t})\sum_{\bm{y}_{>t}}q(\bm{y}_{>t}|c_{\leq t})]\log\frac{q(y_{t}|c_{<t})}{p(y_{t}|c_{<t})}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\qquad+\sum_{t}[\sum_{\bm{y}_{<t}}q(c_{<t})\sum_{y_{t}}q(y_{t}|c_{<t})\sum_{\bm{y}_{>t}}q(\bm{y}_{>t}|c_{\leq t})]\log\frac{q(y_{t}|c_{<t})}{p(\bm{y}_{K,t}|\bm{y}_{K,<t})}$
$=-\sum_{t}\sum_{\bm{y}_{<t}}q(\bm{y}_{<t})\sum_{y_{t}}q(y_{t}|c_{<t})\log\frac{q(y_{t}|c_{<t})}{p(y_{t}|c_{<t})}\sum_{\bm{y}_{>t}}q(\bm{y}_{>t}|c_{\leq t})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\qquad+\sum_{t}\sum_{\bm{y}_{<t}}q(\bm{y}_{<t})\sum_{y_{t}}q(y_{t}|c_{<t})\log\frac{q(y_{t}|c_{<t})}{p(\bm{y}_{K,t}|\bm{y}_{K,<t})}\sum_{\bm{y}_{>t}}q(\bm{y}_{>t}|c_{\leq t}).$
Since $\sum_{\bm{y}_{>t}}q(\bm{y}_{>t}|c_{\leq t})=1$, we can remove $\sum_{\bm{y}_{>t}}q(\bm{y}_{>t}|c_{\leq t})$ in Eq.(31):
$-\sum_{t}\sum_{\bm{y}_{<t}}q(\bm{y}_{<t})\sum_{y_{t}}q(y_{t}|c_{<t})\log\frac{q(y_{t}|c_{<t})}{p(y_{t}|c_{<t})}\sum_{\bm{y}_{>t}}q(\bm{y}_{>t}|c_{\leq t})$ (32)
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\qquad+\sum_{t}\sum_{\bm{y}_{<t}}q(\bm{y}_{<t})\sum_{y_{t}}q(y_{t}|c_{<t})\log\frac{q(y_{t}|c_{<t})}{p(\bm{y}_{K,t}|\bm{y}_{K,<t})}\sum_{\bm{y}_{>t}}q(\bm{y}_{>t}|c_{\leq t})$
$=$ $-\sum_{t}\sum_{\bm{y}_{<t}}q(\bm{y}_{<t})\sum_{y_{t}}q(y_{t}|c_{<t})\log\frac{q(y_{t}|c_{<t})}{p(y_{t}|c_{<t})}+\sum_{t}\sum_{\bm{y}_{<t}}q(\bm{y}_{<t})\sum_{y_{t}}q(y_{t}|c_{<t})\log\frac{q(y_{t}|c_{<t})}{p(\bm{y}_{K,t}|\bm{y}_{K,<t})}$
$=$ $-\sum_{t}\sum_{\bm{y}_{<t}}q(\bm{y}_{<t})[\sum_{y_{t}}q(y_{t}|c_{<t})\log\frac{q(y_{t}|c_{<t})}{p(y_{t}|c_{<t})}-\sum_{y_{t}}q(y_{t}|c_{<t})\log\frac{q(y_{t}|c_{<t})}{p(\bm{y}_{K,t}|\bm{y}_{K,<t})}]$
$=$ $-\sum_{t}\mathbb{E}_{\bm{y}_{<t}\sim q}[\sum_{y_{t}}q(y_{t}|c_{<t})\log\frac{q(y_{t}|c_{<t})}{p(y_{t}|c_{<t})}-\sum_{y_{t}}q(y_{t}|c_{<t})\log\frac{q(y_{t}|c_{<t})}{p(\bm{y}_{K,t}|\bm{y}_{K,<t})}].$
Now we have a step-wise alignment over all sequence from $q$, i.e., $\mathbb{E}_{\bm{y}_{<t}\sim q}$. We aim to align the draft model with the target LLM at every beam search step. Notably, at each beam search step $T$, the sequence lengths are fixed and is independent with $t$. Therefore, we can rewrite Eq.(32) as:
$-\sum_{t}\mathbb{E}_{\bm{y}_{<t}\sim q}[\sum_{y_{t}}q(y_{t}|c_{<t})\log\frac{q(y_{t}|c_{<t})}{p(y_{t}|c_{<t})}-\sum_{y_{t}}q(y_{t}|c_{<t})\log\frac{q(y_{t}|c_{<t})}{p(\bm{y}_{K,t}|\bm{y}_{K,<t})}]$ (33)
$=$ $-\mathbb{E}_{\bm{y}_{\leq T}\sim q}\sum_{t=1}^{T}[\sum_{y_{t}}q(y_{t}|c_{<t})\log\frac{q(y_{t}|c_{<t})}{p(y_{t}|c_{<t})}-\sum_{y_{t}}q(y_{t}|c_{<t})\log\frac{q(y_{t}|c_{<t})}{p(\bm{y}_{K,t}|\bm{y}_{K,<t})}].$
Since we aim to align at every step $T\in\{1,\dots,L\}$, where $L$ is the length of item identifier in LLM-based recommendation, we further normalize the expression by sequence length to prevent different scales on alignment loss across different steps. As such, for every step $T\in\{1,\dots,L\}$, the objective can be rewritten as
$-\mathbb{E}_{\bm{y}_{\leq T}\sim q}\frac{1}{|y_{T}|}\sum_{t=1}^{|y_{T}|}[\sum_{y_{t}}q(y_{t})\log\frac{q(y_{t}|c_{<t})}{p(y_{t}|c_{<t})}-\sum_{y_{t}}q(y_{t})\log\frac{q(y_{t}|c_{<t})}{p(\bm{y}_{K,t}|\bm{y}_{K,<t})}].$ (34)
However, the expectation over the sequence space, (i.e., $\mathbb{E}_{\bm{y}_{\leq T}\sim q}$) is intractable, so we follow previous work (Wen et al., 2023; Kim & Rush, 2016) to approximate it by sampling top-$K$ sequences generated by draft model $\mathcal{M}_{q}$. Now we can rewrite our alignment objective for strict top-$K$ verification and obtain Eq.(3) in our paper:
$\mathop{\arg\max}_{\theta\in\Theta}-\mathbb{E}_{(\bm{x},\mathcal{Y})\sim D^{\prime}}\sum_{\bm{y}\in\mathcal{Y}}\frac{1}{|\bm{y}|}\sum_{t=1}^{\bm{|y|}}[\sum_{y_{t}}q(y_{t}|c_{<t})\log\frac{q(y_{t}|c_{<t})}{p(y_{t}|c_{<t})}-\sum_{y_{t}}q(y_{t}|c_{<t})\log\frac{q(y_{t}|c_{<t})}{p(\bm{y}_{K,t}|\bm{y}_{K,<t})}]$ (35)
$=$ $\mathop{\arg\min}_{\theta\in\Theta}\mathbb{E}_{(\bm{x},\mathcal{Y})\sim D^{\prime}}\sum_{\bm{y}\in\mathcal{Y}}\frac{1}{|\bm{y}|}\sum_{t=1}^{|\bm{y}|}[\sum_{y_{t}}q(y_{t}|c_{<t})\log\frac{q(y_{t}|c_{<t})}{p(y_{t}|c_{<t})}-\sum_{y_{t}}q(y_{t}|c_{<t})\log\frac{q(y_{t}|c_{<t})}{p(\bm{y}_{K,t}|\bm{y}_{K,<t})}],$
where $\mathcal{Y}$ denotes the top-$K$ beam sequences generated from the mixture distribution of draft model and target LLM (see “Alignment Objective” paragraph in Section 3.1).
A.2.3 Derivatives of Acceptance Rate for Relaxed Sampling Verification.
@@
Based on the with-replacement sampling approximation, for every independent sampling, we have acceptance rate of a sequence
@@
Proof.
$b=\mathbb{E}_{\bm{y}\sim q(\bm{y})}\begin{cases}1,&q(\bm{y})\leq p(\bm{y})\\ \frac{p(\bm{y})}{q(\bm{y})},&q(\bm{y})>p(\bm{y})\end{cases}=\mathbb{E}_{y\sim q(\bm{y})}\min(1,\frac{p(\bm{y})}{q(\bm{y})})=\sum_{\bm{y}}\min(1,\frac{p(\bm{y})}{q(\bm{y})})=1-\text{TVD}(p(\bm{y}),q(\bm{y})).$
Based on the with-replacement sampling approximation (see Appendix A.2.1 for proof), we can obtain the acceptance rate of a step with $K$ accepted sequences as $\beta=\prod_{K}b$. Similarly, we could also extend the distribution equivalence for sampling a sequence between the draft model and the target LLM from Appendix A.1 in  (Leviathan et al., 2023). As such, when the sampling-based beam search is performed with a beam size $K$, the joint distribution of the output $K$ sequences from the SD under the relaxed sampling verification is approximately equivalent to the counterparts from the target LLM, according to the approximation of with-replacement sampling, i.e., independent sampling.
@@ A.3 Experimental Details and Analysis
A.3.1 Detailed Experimental Settings
Table 2: Statistics of the datasets.
@@ A.3.1 Detailed Experimental Settings
Recommendation Loss. The recommendation loss used for the backbone LLM-based recommender model in our work is defined as
$\mathcal{L}_{\text{Rec}}=-\frac{1}{N}\sum_{(\bm{x},\bm{y})\sim\mathcal{D}}\sum_{t=1}^{|\bm{y}|}\log\mathcal{M}_{p}({y}_{t}|\bm{y}_{<t},\bm{x}),$ (37)
where $\bm{x}$ is user’s historical interactions, $\bm{y}$ is the user’s next interacted item identifier, and $\mathcal{D}=\{(\bm{x},\bm{y})\}$ denotes the original recommendation dataset. Nevertheless, our method is model-agnostic, thus the recommendation loss $\mathcal{L}_{\text{Rec}}$ can be can be substituted by any form of the loss function from the backend LLM-based recommender models.
Implementation of Tree-based Attention. We first compress the $N$ beam sequences into a single flattened sequence, and then construct the sparse tree-based attention mask for efficient target LLM verification. More precisely, given $\gamma N$ drafted beam sequences with different lengths, where $N$ is the draft beam size and $\gamma$ is the number of drafted beam steps, we first 1) flattened the beam sequences by sequentially adding the newly generated tokens from the beam sequences at each step. We denote the length of the flattened sequence as $L_{f}$. Then, based on the flattened sequence, 2) we construct an attention mask with the shape of $L_{f}\times L_{f}$. Specifically, each row in attention mask represents a specific beam sequence. For each row, we set the corresponding column of the last token as well as the preceding tokens in the beam sequence as 1, otherwise 0. For example, we set the beam size of 2 and collect the beam sequences of 3 steps as shown in Table 3.
step 1 beam 1: a1
step 1 beam 2: a2
step 2 beam 1: a1 b1
step 2 beam 2: a2 b2
step 3 beam 1: a1 b1 c1
step 3 beam 2: a1 b1 c2
Table 3: Beam sequences from 3 beam search steps with beam size 2. The flattened sequence is “a1 a2 b1 b2 c1 c2”.
The flattened sequence will be “a1 a2 b1 b2 c1 c2”. The constructed tree-based attention is shown in Figure 2(c) of our manuscript, where each row represents a specific beam sequence. For each row, the tickled cell represents the preceding tokens and the last tokens of each beam. And the different colors represent the different steps of beam search. This flattened sequence and sparse tree-based attention enable efficient verification since it saves the repeated calculation of the same prefix across different beam sequences, e.g., “a1 b1”.
A.3.2 Additional Experiment Results.
Table 4: Full results of walltime speedup (WS@$K$), accept steps (AS@$K$) between the baselines and AtSpeed instantiated on LC-Rec (LLaMA-7B) across two datasets. For each verification strategy, the best results are highlighted in bold and the second-best results are underlined.
Beauty
DARE 1.13 1.07 1.06 1.15 1.48 0.81 0.44 0.26 0.05 0.00
SFT 1.88 1.55 1.43 1.37 1.55 2.19 1.66 1.32 0.66 0.09
WordKD 2.02 1.70 1.58 1.52 1.58 2.33 1.89 1.60 1.03 0.16
TVDKD 1.84 1.55 1.44 1.37 1.57 2.15 1.67 1.31 0.65 0.09
[…]
AtSpeed-S 2.33 1.97 1.84 1.87 1.84 2.58 2.20 2.00 1.64 0.57
[…]
DARE 1.61 1.65 1.70 1.53 1.95 2.00 2.00 1.97 1.14 1.00
SFT 1.77 1.76 1.80 2.06 2.36 2.15 2.06 2.03 1.99 1.48
WordKD 1.74 1.75 1.81 1.99 2.05 2.11 2.04 2.01 1.87 1.07
TVDKD 1.80 1.77 1.81 2.06 2.35 2.17 2.07 2.03 1.99 1.45
[…]
Games
DARE 0.94 0.95 0.99 1.13 1.44 0.15 0.00 0.00 0.00 0.00
SFT 1.82 1.45 1.43 1.40 1.58 2.13 1.49 1.32 0.91 0.15
WordKD 1.82 1.45 1.31 1.35 1.47 2.13 1.49 1.10 0.80 0.14
TVDKD 1.70 1.33 1.24 1.32 1.50 1.99 1.26 0.95 0.66 0.09
[…]
DARE 1.59 1.64 1.68 1.19 1.42 2.00 2.00 1.96 0.37 0.00
SFT 1.96 1.85 1.84 1.97 1.69 2.28 2.12 2.05 1.89 0.58
WordKD 1.84 1.77 1.78 1.84 1.56 2.17 2.05 1.99 1.68 0.25
TVDKD 1.96 1.81 1.81 1.90 1.55 2.26 2.08 2.02 1.80 0.29
SeqKD 1.91 1.87 1.90 2.03 2.05 2.22 2.13 2.10 1.98 1.22
[…]
Comprehensive Results on Beauty and Games Datasets. We report the full acceleration performance comparison on Beauty and Games datasets in Table 4 and the full recommendation performance comparison on the two datasets in Table 5, where we can have similar observations as in Table 1.
Table 5: Full results of ranking performance in terms of Recall and NDCG between the baselines and AtSpeed instantiated on LC-Rec (LLaMA-7B) on Beauty and Games datasets.
Beauty
Method Recall@5 Recall@10 NDCG@5 NDCG@10
[…]
Target LLM (Sampling) 0.0056 0.0082 0.0043 0.0066
Relaxed Sampling Verification DARE 0.0059 0.0102 0.0044 0.0060
SFT 0.0057 0.0091 0.0041 0.0063
WordKD 0.0066 0.0105 0.0043 0.0058
TVDKD 0.0057 0.0083 0.0045 0.0054
SeqKD 0.0055 0.0116 0.0045 0.0067
AtSpeed-S 0.0060 0.0096 0.0046 0.0060
AtSpeed-R 0.0058 0.0092 0.0049 0.0063
Average 0.0059 0.0098 0.0045 0.0061
Games
Method Recall@5 Recall@10 NDCG@5 NDCG@10
[…]
Target LLM (Sampling) 0.0075 0.0115 0.0066 0.0079
Relaxed Sampling Verification DARE 0.0076 0.0119 0.0065 0.0080
SFT 0.0073 0.0112 0.0060 0.0074
WordKD 0.0072 0.0113 0.0058 0.0073
TVDKD 0.0069 0.0108 0.0061 0.0074
SeqKD 0.0071 0.0110 0.0059 0.0073
AtSpeed-S 0.0080 0.0131 0.0068 0.0085
AtSpeed-R 0.0076 0.0123 0.0063 0.0080
Average 0.0074 0.0117 0.0062 0.0077
[…]
Table 6: Performance comparison between AtSpeed and baselines on MovieLens-1M dataset. For each verification strategy, the best results are highlighted in bold and the second-best results are underlined.
MovieLens-1M
Verification Method WS@3 WS@5 WS@10 WS@20 AS@3 AS@5 AS@10 AS@20
Strict Top-$\bm{K}$ DARE 1.26 1.28 1.39 1.72 1.00 1.00 1.00 1.00
SFT 1.65 1.35 1.48 1.76 1.99 1.26 1.14 1.03
WordKD 1.29 1.29 1.39 1.69 1.07 1.02 1.00 1.00
TVDKD 1.73 1.72 1.25 1.24 1.98 1.90 1.04 1.00
[…]
AtSpeed-S 1.86 1.79 1.80 1.75 2.08 2.03 1.98 1.09
[…]
Relaxed Sampling DARE 2.01 1.84 1.35 1.44 2.16 2.02 1.00 0.35
SFT 2.08 2.03 2.02 1.61 2.28 2.20 2.06 0.93
WordKD 1.97 1.87 1.48 1.28 2.15 2.08 1.23 0.00
TVDKD 2.00 1.98 1.68 1.29 2.30 2.21 1.59 0.00
SeqKD 2.13 2.08 1.93 1.56 2.19 2.14 2.01 0.78
[…]
Goodreads
Verification Method WS@3 WS@5 WS@10 WS@20 AS@3 AS@5 AS@10 AS@20
Strict Top-$\bm{K}$ DARE 1.30 1.32 1.44 1.75 1.00 1.00 1.00 1.00
SFT 1.83 1.81 2.17 2.46 2.04 1.98 1.72 1.07
WordKD 1.83 1.92 2.07 2.38 2.00 1.96 1.58 1.00
TVDKD 1.89 1.93 2.17 2.46 2.07 1.97 1.70 1.07
SeqKD 1.82 1.89 2.19 2.48 2.00 1.96 1.73 1.08
[…]
Relaxed Sampling DARE 1.84 1.83 1.35 1.43 2.06 2.02 1.00 0.35
SFT 2.15 2.09 1.70 1.91 2.27 2.08 1.01 0.10
WordKD 2.01 2.04 1.68 1.92 2.15 2.05 1.00 0.15
[…]
SeqKD 1.90 1.96 1.66 1.85 2.08 2.01 1.00 0.02
[…]
AtSpeed-R 2.45 2.39 1.77 2.36 2.50 2.32 1.10 0.87
[…]
@@
Utilization of Beam Sequences from All Steps. Table 7 shows inference and training efficiency comparison between AtSpeed training with and without additional data generated from beam search results on diverse steps. From the table, it is observed that incorporating data from diverse generating steps usually enhances the speedup and accept steps, which is attributed to the additional alignment of top-$K$ sequences across different steps. Nevertheless, using additional data could incur increased training time and storage cost. Specifically, utilizing a larger dataset necessitates greater storage capacity and extends the training duration. Despite these additional costs, the marginal improvement in performance is not substantial, leading to a low cost-effectiveness ratio. Therefore, we should carefully consider the balance between the benefits of enhanced alignment against the increased training resource burden.
@@
Hyper-parameter Analysis. 1) Effect of alignment strength $\alpha$. Figure 5(b-c) presents the results indicating the impact of alignment strength on AtSpeed-R. As the value of $\alpha$ increases, the overall trend of accept steps exhibits an upward trajectory from $0.1$ to $0.9$, which validates the necessity of alignment to improve the performance. Nevertheless, compared to AtSpeed-S (Figure 4(b-c)), we can find that only extremely large $\alpha$ (e.g., 1) would hurt the performance of AtSpeed-R while setting $\alpha$ to 0.7 for AtSpeed-S already cause a performance drop. We suspect the different behaviors are due to the different scales between $L_{\text{Align-S}}$ and $L_{\text{Align-R}}$. Specifically, given the sequence distribution from draft model $q$ and that from target LLM $p$, we have RKLD $q\log\frac{q}{p}$ in AtSpeed-S and TVD $\frac{|q-p|}{2}$ in AtSpeed-R to align the two models. When sequence probability $q$ becomes small as the sequence length increases, $\frac{q}{p}$ becomes very large. It is possible for AtSpeed-S to give a larger loss value compared to AtSpeed-R (i.e., $\frac{|p-q|}{2}$). Therefore, for the same $\alpha<1$, the strength that is too large for AtSpeed-S might still benefit alignment for AtSpeed-R.
2) Effect of draft length $\gamma$. Figure 6 presents the impact of varying $\gamma$ on the performance of AtSpeed. Despite acceptance generally increases with $\gamma$, speedup does not consistently improve with $\gamma$. The trend of speedup in relation to $\gamma$ is influenced by draft beam size parameter $N$. Specifically, when $N$ is small, speedup tends to increase with $\gamma$; conversely, when $N$ is large, speedup tends to decrease with $\gamma$. This variation is primarily attributed to the computational overhead during the draft phase, which escalates with larger beam size, thereby diminishing the acceleration benefits. The key determinant of the draft phase’s computational cost is the constraint search process.